40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Tue Jul 2 11:21:24 2013, Erica, DailyProgress, fiber optic, mode matching snip_of_beam_profiling_6_27.JPGP1020518.JPG
    Reply  Wed Jul 3 09:46:48 2013, Erica, DailyProgress, fiber optic, mode matching 130703-095332.jpg
       Reply  Thu Jul 11 01:40:19 2013, Erica, DailyProgress, fiber optic, mode matching P1020690.JPGP1020692.JPG
          Reply  Fri Jul 12 00:51:39 2013, Erica, Notes, fiber optic, mode matching P1020705.JPGP1020703.JPG7_11_13_beam_profile.PNG7_11_13_beam_from_fixed_lens.PNG
             Reply  Tue Jul 16 10:27:47 2013, Erica, Notes, fiber optic, mode matching, beam recombination 
                Reply  Tue Jul 16 11:02:47 2013, Erica, Notes, fiber optic, mode matching, beam recombination 130805-120628.jpg
                   Reply  Wed Jul 17 11:58:47 2013, Erica, Notes, fiber optic, mode matching, beam recombination 7_16_13_determining_polarization_axis.PNGTEK0000.PNG
                      Reply  Wed Jul 17 18:29:37 2013, Erica, DailyProgress, fiber optic, reducing drift in the recombined beam 
                         Reply  Sat Jul 20 23:39:09 2013, Erica, DailyProgress, fiber optic, reducing drift in the recombined beam 
                            Reply  Tue Jul 23 10:29:12 2013, Erica, DailyProgress, fiber optic, reducing drift in the recombined beam shielded_10s_TEK0005.PNGunshielded_10s_TEK0027.PNG
                               Reply  Tue Jul 23 11:43:00 2013, Evan, DailyProgress, Electronics Equipment, Zach's op-amp recommendations 
                               Reply  Fri Jul 26 11:07:44 2013, Erica, DailyProgress, BEAT, circuit for measuring temperature fluctations on CTN table P1030056.JPGP1030057.JPGP1030059.JPGP1030056.JPG
                                  Reply  Fri Jul 26 11:22:45 2013, Erica, DailyProgress, BEAT, circuit for measuring temperature fluctations on CTN table 
                                     Reply  Fri Jul 26 11:45:28 2013, Erica, DailyProgress, BEAT, circuit for measuring temperature fluctations on CTN table 
                                        Reply  Mon Jul 29 13:30:26 2013, Erica, DailyProgress, BEAT, adding resistor, capacitor, and sockets w_o_capacitor_7_26_13.JPGcapacitor_7_26_13.JPGP1030066.JPGP1030070.JPG
                                           Reply  Tue Jul 30 12:23:41 2013, Erica, DailyProgress, BEAT, addressed noise problem, DC signal 
                                              Reply  Thu Aug 1 00:25:02 2013, Erica, DailyProgress, fiber optic, insulating foam 731_7graph.PNG731_11graph.PNGP1030318.JPG
                                                 Reply  Fri Aug 2 12:38:41 2013, tara, DailyProgress, fiber optic, insulating foam chmbrtmp.JPGfringe.JPG
                                                    Reply  Tue Aug 6 12:31:24 2013, Erica, DailyProgress, fiber optic, insulating foam P1030403.JPG
                                                       Reply  Wed Aug 7 13:27:00 2013, Erica, DailyProgress, fiber optic, more foam, calculations for convertion from voltage to frequency P1030409.JPGP1030417.JPGP1030420.JPG
                                                          Reply  Wed Aug 7 18:53:20 2013, Erica, DailyProgress, fiber optic, noise budget, Matlab notes. noise_budget.PNGP1030427.JPG
                                                             Reply  Tue Aug 13 12:39:36 2013, Erica, DailyProgress, fiber optic, rewrote calculations for converting from voltage noise to frequency noise P1030458.JPGP1030459.JPGP1030460.JPG
                                                       Reply  Tue Aug 13 12:28:20 2013, Erica, Notes, Electronics Equipment, new AD 590 circuit and its noise calculation, beam alignment P1030461.JPGP1030454.JPGP1030456.JPGP1030457.JPGAD590circuit2.PNG
                                                          Reply  Fri Aug 16 10:35:44 2013, Erica, Notes, Electronics Equipment, AD590 msrmt, conversion to freq noise 
                   Reply  Tue Jul 30 16:19:32 2013, Erica, Notes, fiber optic, setup for beam recombination new_setup.PNGbeam_recombination2.PNG
                   Reply  Mon Aug 5 12:21:00 2013, Erica, Notes, fiber optic, calculation: phase change for fiber 130805-121449.jpg
    Reply  Tue Jul 30 15:48:47 2013, Erica, DailyProgress, fiber optic, noise for GYRO 
Message ID: 1295     Entry time: Tue Aug 13 12:28:20 2013     In reply to: 1282     Reply to this: 1301
Author: Erica 
Type: Notes 
Category: Electronics Equipment 
Subject: new AD 590 circuit and its noise calculation, beam alignment 
Attached are noise calculations for a new circuit for the AD590.

The largest noise source was the AD 590, as desired, with a noise of 8*10^-6 V/rtHz. We wanted this to be the dominant source of noise.
The largest the other components had was 6*10^-7 V/rtHz, from R1.

Originally, we were going to use a potentiometer where for R2, but Rana said that the potentiometer would be unstable; the resistance would change in response to temperature changes. He said we should play around using different fixed resistors. We ended up using the potentiometer to determine approximately what R2 should be and when doing so, I did notice that holding the potentiometer made the resistance fluctuate.



I could put a 1M ohm resistor for R3 and a 27 kOhm one for R4 for a larger voltage readout

I finished soldering it together.


Tara and Evan put in some half-wave and quarter wave plates in the beam path, which misaligned the beam.
However, Evan and I could not realign the beam to get an output of more than 5 uW coming from 2mW after trying for a long time.
Attachment 1: P1030461.JPG  3.462 MB  | Show | Hide all | Show all
Attachment 2: P1030454.JPG  3.444 MB  | Hide | Hide all | Show all
P1030454.JPG
Attachment 3: P1030456.JPG  3.301 MB  | Hide | Hide all | Show all
P1030456.JPG
Attachment 4: P1030457.JPG  3.474 MB  | Hide | Hide all | Show all
P1030457.JPG
ELOG V3.1.3-